OpenStudy (anonymous):

find the integral (4x^3+2/2x^4+4x+3) dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{4x ^{3}+2}{2x ^{4}+4x+3}dx\]

OpenStudy (anonymous):

I'm thinking a u-sub of u=2x^{4}+4x+3 will work, right?

OpenStudy (anonymous):

my answer choices are \[2 \ln \left| 2x ^{4}+4x+3 \right|+C\] , \[\frac{-2}{(2x ^{4}+4x+3)^{2}}+C\] , \[1/2\ln \left| 2x ^{4}+4x+3 \right|+C\] , \[\frac{-1}{2(2x ^{4}+4x+3)^{2}}+C\]

OpenStudy (anonymous):

we gotta try partial fraction

OpenStudy (anonymous):

okay where do we go from there

OpenStudy (anonymous):

It's actually a simple U-sub

OpenStudy (anonymous):

its 3rd choice do you need to show work?

OpenStudy (anonymous):

let u=2x^4 +4x + 3 evalute du, du = 8x^3 + 4 = 1/2 (4x^3 +2) then the rest is a general natural log integral

OpenStudy (anonymous):

are you sure it is the third choice? no i just need the correct answer

OpenStudy (anonymous):

yep i used u substitution just like wattagwan

OpenStudy (anonymous):

sweet thank you so much

OpenStudy (anonymous):

no problem